OpenStudy (mahogany):

A 8.6 kg block initially at rest is pulled to the right along a horizontal, frictionless surface by a constant, horizontal force of 16.3 N. Find the speed of the block after it has moved 4.3 m. Answer in units of m/s.

OpenStudy (matt101):

In the future you should post this in the physics section! You're given mass (m), force (F), distance (x), and initial speed (vi), and asked to find final speed (vf). If we can figure out some equations that use these variables only, we can solve for the unknown. You know from Newton's Second Law that F = ma, or a = F/m. You also know from the equations of linear motion that vf^2 = vi^2 + 2ax, or a = (vf^2-vi^2)/(2x). We have isolated a in both equations, and so can set them equal to each other to get F/m = (vf^2-vi^2)/(2x). Solve that equation for vf and you will have your answer!